282006:

URGENT: The seventh-grade class is building target areas for a PE activity. The bases for the game will be circular in shape. The diameter of each circle is 5 feet. Approximately how many square feet of the turf need to be painted for a base circle? Use 3.14 for π and round your answer to the nearest tenth. @umm @TheSmartOne

umm:

Is there answer choices to this, or self answered?

282006:

there are choices 15.7 square feet 19.6 square feet 42.8 square feet 78.6 square feet

umm:

Okay. Do you know what the area of the circle would be? I'm trying to break it down for you to understand, since knowingly direct answers aren't allowed.

282006:

are yes

282006:

i ment i do know what it is

umm:

Okay - so the formula is A = πr2 (pi times squared r). So, in other words, you cut the diameter in half (5 ÷ 2 = 2.5 feet) and do this: 2.5 × 2.5 (this is the squared 2.5).

282006:

2.5×2.5=6.25

282006:

6.25×3.14=19.625

umm:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @282006 6.25×3.14=19.625 \(\color{#0cbb34}{\text{End of Quote}}\) Which makes the answer: 19.6 (choice B.)

282006:

thank you

umm:

You're welcome.